# ACPO: AI-Enabled Compiler-Driven Program Optimization

### v2

@@
Abstract.
The key to performance optimization of a program is to decide correctly when a certain transformation should be applied by a compiler. This is an ideal opportunity to apply machine-learning models to speed up the tuning process; while this realization has been around since the late 90s, only recent advancements in ML enabled a practical application of ML to compilers as an end-to-end framework.
@@ Abstract.
This paper presents ACPO: AI-Enabled Compiler-driven Program Optimization; a novel framework to provide LLVM with simple and comprehensive tools to benefit from employing ML models for different optimization passes. We first showcase the high-level view, class hierarchy, and functionalities of ACPO and subsequently, demonstrate a couple of use cases of ACPO by ML-enabling the Loop Unroll and Function Inlining passes and describe how ACPO can be leveraged to optimize other passes. Experimental results reveal that ACPO model for Loop Unroll is able to gain on average 4% compared to LLVM’s O3 optimization when deployed on Polybench. Furthermore, by adding the Inliner model as well, ACPO is able to provide up to 4.5% and 2.4% on Polybench and Cbench compared with LLVM’s O3 optimization, respectively.
30(2,0.5)
LLVM Dev 2023 Presentation: https://youtu.be/jzT-tar_X0U
1. Introduction
@@
State-of-the-art compilers are built with sophisticated optimization pipelines that result in robust, reproducible (Ivie and Thain, 2018), and reliable code generation (Aho et al., 1986; Hall et al., 2009). The LLVM compiler framework is no exception with its multi-layer design, where passes in the mid-end can perform analyses or transformations on a code region, one after another (Lattner, 2008). LLVM passes operate on code regions with varying levels of granularity, including loops, functions, call graphs, and modules. Compiler engineers have continuously improved the quality of the optimization passes in LLVM, and have derived a set of fixed-length optimization pipelines, known as standard optimization levels, i.e., O{1,2,3,s,z}, which are designed to improve the code size or performance (to varying degree) of the programs to which they are applied. At the highest optimization level available, i.e. O3, more than 60 unique transformation passes exist in LLVM’s pipeline (Ashouri et al., 2017).
@@
Leveraging machine learning (ML) for compilers has become an extensively researched topic in the past years. In many instances, the authors showed experimental results that outperformed LLVM’s O3 (Agakov et al., 2006; Chen et al., 2012; Park et al., 2013; Ashouri et al., 2016, 2017; Cummins et al., 2017; Georgiou et al., 2018; Haj-Ali et al., 2020; Ashouri et al., 2022) or Deep Neural Networks (DNN) kernel schedulers (Patabandi et al., 2021; Patabandi and Hall, 2023) and lately, using Large Language Models (LLMs) (Cummins et al., 2023; Gong et al., 2024). Machine learning approaches normally rely on data generated by an autotuner (Ansel et al., 2014; Kalyan et al., 2019), which helps with exploring an optimization space formed by the number of parameters in the decision set. Alternatively, by the growth of applications of Reinforcement Learning (RL), as opposed to generating offline training data by an autotuner, authors have explored their action space, online, and during the training process of the RL agent. This is done by means of feeding back a reward value to update the agent’s policy (Trofin et al., 2021; Cummins et al., 2021b; Ashouri et al., 2022; Wang et al., 2022a). To characterize the software being compiled for the ML model, program representation methods are employed, which generally fall into three categories: using static features (Fursin et al., 2011; Ashouri et al., 2022), using dynamic features (Cavazos et al., 2006; Hoste and Eeckhout, 2007), and graph-based characterization (Park et al., 2012; Ben-Nun et al., 2018; Alon et al., 2019; Cummins et al., 2021a). In recent years, there have been considerable advancements in the space, especially in the graph-based program representation techniques (Guo et al., 2020; Cummins et al., 2021a) and token-based identifiers (Wang et al., 2021), which are outside the scope of this work.
The vision for ACPO is to enable the use of Machine Learning in a compiler, in a way that is modular, extendable, and easy to use. ACPO aims to allow for easy collaboration between compiler developers and ML architects by use of clear and simple APIs that enable the collection of relevant data for training purposes, providing inputs to an ML model during inference, and separating the compiler from the model and ML framework to enable model development and testing within the compiler context with minimal intrusion into the compiler code base. This enables the compiler engineers to optimize passes with an easy-to-use set of tools and limited ML knowledge. At the same time, once an ML model is integrated into the compiler, it can be easily optimized and redesigned with no changes to the compiler code base, effectively enabling ML architects to contribute to compiler improvement. This is possible due to ACPO’s separation of the ML framework and the compiler, making it possible for the compiler to work just as easily with any ML model as it is to work with any ML framework, e.g., TensorFlow (Abadi et al., 2016), PyTorch (Paszke et al., 2017), etc, and even other representation mechanism (Fursin et al., 2008; Hoste and Eeckhout, 2007; Cummins et al., 2021a). To this end, ACPO proposes the following contributions:
@@
As a framework, it provides a comprehensive set of tools, libraries, and algorithmic methods by enabling compiler engineers with a straightforward and user-friendly interface to instantiate ACPOModel classes to replace LLVM’s existing hard-coded heuristics.
@@
ML APIs and the compiler are seamlessly connected, but at the same time, they are not interdependent; changing compiler versions, ML models, or ML frameworks won’t break the functionality of ACPO. As long as the inputs/outputs match, users can easily revise the ML side without the need to rebuild LLVM after every change.
@@
Showcasing the benefits of ACPO, we demonstrate two different scenarios:
Loop Unroll Pass — Building both the interface and the ML model. Loop Unroll pass can be optimized from scratch by adding the necessary code to instantiate ACPOLUModel from ACPOModel and also by providing an ML model to outperform LLVM’s unrolling decisions at O3 on Polybench and a number of other benchmarks.
Function Inlining Pass — Building the interface and leveraging an existing ML model. Function Inlining pass can be optimized by leveraging an existing ML model, i.e., MLGOPerf (Ashouri et al., 2022), using ACPO framework. We do that by creating an ACPOModel instance, called ACPOFIModel, and ensuring its inputs and outputs match the existing model so it can be simply invoked at the appropriate point in the compilation process. We demonstrate the ease of integration and versatility of ACPO framework and that how it can be used to connect existing models to the new frameworks. Experimental results on Cbench reveal that the performance of ACPOFIModel is comparable with the reported results in the original work while benefiting from a multi-layer design of ACPO on the separation between the ML APIs and the compiler.
@@
The rest of the paper is organized as follows. Section 2 discusses the state of the art. Section 3 gives details on our proposed framework. Under Sections 4.1 and 4.2, we showcase how Loop Unroll and Function Inlining passes can benefit by instantiating ACPOLUModel. Section 5 showcases the experimental results including the individual and the combined results of deploying both models and Finally, in Section 6 we discuss challenges and propose directions for future work.
2. Related Work
@@
A recent survey on the compiler autotuning and optimization (Ashouri et al., 2018) reveals that there have been a vast number of approaches to employ ML techniques for automatically constructing optimization heuristics. These approaches mainly focused on two major problems of compiler optimization: (1) the optimization selection (Bodin et al., 1998; Cooper et al., 2002; Agakov et al., 2006; Fursin and Temam, 2009; Park et al., 2013; Ashouri et al., 2016) and (2) the phase ordering problem (Kulkarni and Cavazos, 2012; Ashouri et al., 2017; Nobre et al., 2018; Huang et al., 2019; Mammadli et al., 2020; Cummins et al., 2021b; Wang et al., 2022a). Further classification of the two aforementioned problems and by the type of the approach, provides insights for categorizing the works into three groups: (1) autotuning-based approaches (Fursin et al., 2011; Chen et al., 2012; Ashouri et al., 2013; Ansel et al., 2014; Kalyan et al., 2019; Cummins et al., 2021b; Wang et al., 2022b; Park et al., 2022), (2) end-to-end frameworks of optimizations (Fursin et al., 2008; Cummins et al., 2017; Haj-Ali et al., 2020; Cummins et al., 2021b; Rotem and Cummins, 2021; Das et al., 2020; Mannarswamy and Das, 2022; Ashouri et al., 2022; Wang et al., 2022b), and (3) ML-guided approaches (Rotem and Cummins, 2021; Trofin et al., 2021; Ashouri et al., 2022).
@@
Autotuning-based approaches leverage an iterative methodology to wrap around the compiler and to speed up the search in finding better solutions to optimizing a single or multiple passes, often constructing a new sequence of optimization to outperform standard optimization levels by means of phase-ordering or optimization selection (Agakov et al., 2006; Ashouri et al., 2016; Cummins et al., 2021b; Wang et al., 2022a). This can be done in a coarse (Kulkarni and Cavazos, 2012; Ashouri et al., 2016, 2017) or fine-grained manner (Kalyan et al., 2019; Phothilimthana et al., 2021; Cummins et al., 2021b). End-to-end approaches often provide a framework that can be paired or built with the compiler and to offer boosted search or ordering of optimizations (Cummins et al., 2017; Haj-Ali et al., 2020; Mannarswamy and Das, 2022). ML-guided approaches build the decision-making capabilities of ML models into single or multiple passes that can act as a standalone ML-enabled pass. Their corresponding ML models can be either built with the compiler using an Ahead-of-Time (AOT) compilation or be used at runtime using an alternative interprocess communication method, e.g., Python named pipes. All things considered, several recent works can be included in multiple classes, therefore, the lists here are not mutually exclusive.
@@
Cummins et al. (Cummins et al., 2021b) propose CompilerGym, a fine-grained RL environment to speed up the search for tuning a single or multiple compiler passes defined as action space. The authors leverage OpenAI’s Gym (Brockman et al., 2016) library to provide a flexible environment to tackle the selection and the phase-ordering of optimization. The results showcase the benefits of using their RL agent to further reduce the code size of Cbench benchmarks. Contrary to this work, we are proposing an ML-guided approach to provide a framework to optimize passes. Additionally, our objective is to improve the performance as opposed to reducing the code size of a binary.
@@
Trofin et al. (Trofin et al., 2021) propose MLGO that uses ML to advise the Inliner as to whether or not a call site should be inlined, in order to minimize the size of the generated code. MLGO framework compiles the ML model into C++ libraries using Ahead-of-Time (AOT) compilation which becomes part of the compiler at build time. Additionally, the authors have extended the code base to provide optimization for the register allocation problem. Later, MLGOPerf (Ashouri et al., 2022), extended MLGO to optimize inlining for performance as opposed to code size and showed efficacy on a number of benchmarks, e.g. SPEC CPU 2006 and Cbench.
In this work, we focus on the hard problem of ML performance optimization, where the key challenge is the development of model architecture and the features used to effectively capture key performance use cases and to facilitate the generation of robust ML models. In the process, we develop a scalable, extendible, end-to-end ML-guided framework to enable easy use of ML within optimization passes in compilers such as LLVM. We demonstrate the effectiveness and the generality of our framework by applying ML-enabling Loop Unroll and Function Inlining passes. We show how ML can easily and effectively be incorporated into a compiler while maintaining separation from the underlying ML frameworks. This framework is easily applicable to other passes and tasks within the compiler domain.
3. Proposed Methodology
Figure 1 presents the high-level design of our infrastructure. There are multiple components involved with the ACPO architecture: Autotuner, compiler, ACPOModel abstraction, low/high-level ML APIs, and ML Framework. For the most part, the autotuner is a wrapper around the compiler and its hooks in the compiler are used to generate model training data. During the inference flow, i.e., when ML models are used for decision making within the compiler, the autotuner is not used and we can think of the compiler as being at the top of the stack.
@@ 3. Proposed Methodology
The ACPO inference flow begins when an ML-enabled pass is invoked and creates an instance of ACPOModel. The instance of ACPOModel aggregates appropriate features from the code region being analyzed and specifies the kind of output required from the ML model. The ACPOModel object then transfers the features, along with information about the model being invoked, to our ML framework via a set of remote procedure call APIs. The APIs provide the ability to send requests, including loading an appropriate model into the ML framework, executing an inference call, and returning the result of the inference back to the compiler. The compiler then applies transformations to the input program as prescribed in the inference output it receives.
In the following sections, we describe in more detail the training and inference flows, and then elaborate on each component of the ACPO framework to show how they fit with the training and inference flows.
@@
The autotuner parameters can be either coarse- or fine-grained. The former includes the selection and the phase-ordering of different optimization passes as a black box, as well as optimization parameters that apply to the whole translation unit, while the latter facilitates the tuning of optimization parameters for individual code regions of a program, e.g. instructions, loops, functions, etc., independently of each other. In this work, we perform fine-grained autotuning with an autotuner (Kalyan et al., 2019), which is built on top of OpenTuner (Ansel et al., 2014). This allows us to generate meaningful training data and target labels that capture the effect of optimization parameters on code region types of interest.
3.1. ACPO Training
A key aspect of creating an ML-driven compiler is the training of models on sample data to ensure that the compiler makes decisions consistently with good results contained in the training data. To do this, we opted for a two-step approach: generate data based on a set of benchmarks the model is expected to reasonably reproduce and then train the model with the collected data. The data generation step is shown in Figure 2a.
To generate data for training purposes we rely on the autotuner. To do this efficiently, we enabled what we refer to as fine-grained autotuning to control each decision of a given transformation pass independently of one another. For example, in the case of loop unrolling, we enabled the autotuner to search for a good unroll factor for each loop, instead of applying an unroll threshold that would be specified on the compiler command line and hence applied globally on all loops. With the help of fine-grained autotuning, we are able to specify which pass we wish to tune and apply this kind of tuning to a set of benchmarks. At the compilation stage, with the instructions from the autotuner, the compiler proceeds to optimize an application and eventually reaches the ML-enabled pass for which the autotuner is currently providing decisions.Note that ACPO supports leveraging different scopes of features as long as it is supported by the compiler and it is up to the developer to decide which ACPO pass might benefit from which scope during the process of design and implementation. The features are passed with std::vector of pairs to the ML side. This process can also be scaled to an arbitrary number of features and feature types. The transformation decision as prescribed by the autotuner is applied to the code region (e.g. the autotuner can set the unroll factor of a loop to a specific value), and then the compilation continues normally until it reaches another code region that requires a decision from the ML model, at which time the process is repeated. Once the compilation is completed, the autotuner runs the program and collects a measurement for performance feedback (the simplest metric being the execution time of the program). We use the O3 compilation as our baseline and measure impact in terms of speedup relative to the O3 compilation.
3.2. ACPO Inference

### v3

# ACPO: An AI-Enabled Compiler Framework

@@
Abstract
@@
The key to the performance optimization of a program is to decide correctly when a compiler should apply a certain transformation. This paper presents ACPO: An AI-Enabled Compiler Framework, a novel framework that provides LLVM with simple and comprehensive tools to benefit from employing ML models for different optimization passes. We first showcase the high-level view, class hierarchy, and functionalities of ACPO and subsequently, demonstrate a couple of use cases of ACPO by ML-enabling the Loop Unroll and Function Inlining passes used in LLVM’s O3. and finally, describe how ACPO can be leveraged to optimize other passes. Experimental results reveal that the ACPO model for Loop Unroll can gain on average 4%, 3%, 5.4%, and 0.2% compared to LLVM’s vanilla O3 optimization when deployed on Polybench, Coral-2, CoreMark, and Graph-500, respectively. Furthermore, by including both Function Inlining and Loop Unroll models, ACPO can provide a combined speedup of 4.5% on Polybench and 2.4% on Cbench when compared with LLVM’s O3, respectively.
20(1.6,0.5)
* The short version is accepted at ACM/IEEE CASES’24 – https://ieeexplore.ieee.org/document/10740757
* LLVM-Dev 2023 Presentation: https://youtu.be/jzT-tar_X0U
I Introduction & Related Works
@@
State-of-the-art compilers, e.g., LLVM, are built with sophisticated optimization pipelines that result in robust, reproducible [1], and reliable code generation [2]. The LLVM compiler framework is no exception with its multi-layer design, where passes in the mid-end can perform analyses or transformations on a code region, one after another [3]. LLVM passes operate on code regions with varying levels of granularity, including loops, functions, call graphs, and modules. Compiler engineers have continuously improved the quality of the optimization passes in LLVM, and have derived a set of fixed-length optimization pipelines, known as standard optimization levels, i.e., O{1,2,3,s,z}, which are designed to improve the code size or performance (to varying degrees) of the programs to which they are applied. At the highest optimization level available, i.e. O3, more than 60 unique transformation passes exist in LLVM’s pipeline [4].
@@
Leveraging machine learning (ML) for compilers has recently become an extensively researched topic. In many instances, the authors showed experimental results that outperformed LLVM’s O3 [5, 4, 6, 7, 8, 9, 10, 11, 12]. Machine learning approaches normally rely on data generated by an autotuner [13, 14], which helps explore an optimization space formed by the number of parameters in the decision set. Alternatively, by the growth of applications of Reinforcement Learning (RL), as opposed to generating offline training data by an autotuner, authors have explored their action space, online, and during the training process of the RL agent. This is done by feeding back a reward value to update the agent’s policy [15, 16, 8, 17]. To characterize the software being compiled for the ML model, program representation methods are employed, which generally fall into three categories: using static features [18, 8], using dynamic features [19, 20], and graph-based characterization [21, 22, 23, 24]. In recent years, there have been considerable advancements in the space, especially in the graph-based program representation techniques [25, 24] and token-based identifiers [26], which are outside the scope of this work. Recently, there have been a few works providing an end-to-end approach to construct ML-based heuristics for optimizing the decision-making process of the transformation passes [6, 27, 28, 15, 16, 29, 8]. Among these, only a subset provides an ML-guided optimization approach (MLGO) [16, 28, 15, 8] baked into a single or multiple passes to streamline the optimization process inside the pass functionality readily. The recent advancements in this area beg the question of how scalable, user-friendly, and comprehensive a framework should be to provide seamless integration with existing LLVM optimization passes while maintaining a clean separation of concerns between the ML and the LLVM layers.
MLGO uses ML to advise the Inliner as to whether or not a call site should be inlined to minimize the size of the generated code. Later, MLGOPerf [8] extended MLGO to optimize inlining for performance as opposed to code size and showed efficacy on several benchmarks, e.g. SPEC CPU 2006 and Cbench. In this work, similar to MLGOPerf, we also focus on the hard problem of ML performance optimization, where the key challenge is the development of model architecture and the features used to capture key performance use cases. However, we develop a scalable, extendible, end-to-end ML-guided framework to enable easy use of ML within optimization passes in compilers such as LLVM. Contrary to MLGO, ACPO also provides a standalone and modular feature collection class defined in different scopes of LLVM, i.e., Module, Function, CallGraph, and Loops. Users can easily add selected features to their ACPO passes by the feature names or scopes. To summarize, ACPO proposes the following contributions:
1.
@@
A framework to provide a comprehensive set of predefined program features, libraries, and algorithmic methods by enabling compiler engineers with a user-friendly interface to instantiate ACPOModel classes to replace LLVM’s existing hard-coded heuristics.
2.
@@
ML APIs and the compiler are seamlessly connected, but at the same time, they are not interdependent; changing compiler versions, ML models, or ML frameworks won’t break the functionality of ACPO. As long as the inputs/outputs match, users can easily revise the ML side without rebuilding LLVM after every change.
3.
@@
Showcasing the benefits of ACPO, we demonstrate two different scenarios with LLVM:
a) Loop Unroll Pass — Building both the interface and the ML model. b) Function Inlining Pass — Building the interface and leveraging an existing ML model [8].
@@
The rest of the paper is organized as follows. Section II discusses the state of the art. Section III gives details on our proposed framework and how the different layers of abstraction communicate with each other. Under Sections IV-A and IV-B, we showcase how Loop Unroll and Function Inlining passes can benefit by instantiating ACPOLUModel. Section V showcases the experimental results including the individual and the combined results of deploying both models and Finally, in Section VI we discuss challenges and propose directions for future work.
II Related Work
@@
A survey on the compiler autotuning and optimization [30] reveals that there have been a vast number of approaches to employ ML techniques for automatically constructing optimization heuristics. These approaches mainly focused on two major problems of compiler optimization: (1) the optimization selection [31, 32, 33, 34, 35, 5] and (2) the phase ordering problem [36, 4, 37, 38, 39, 16, 17]. Further classification of the two aforementioned problems and by the type of the approach provides insights for categorizing the works into three groups: (1) autotuning-based approaches [18, 40, 41, 13, 14, 16, 42, 43], (2) end-to-end frameworks of optimizations [44, 6, 7, 16, 28, 27, 29, 8, 42], and (3) ML-guided approaches [28, 15, 8].
@@
Autotuning-based approaches leverage an iterative methodology to wrap around the compiler and to speed up the search in finding better solutions to optimizing a single or multiple passes, often constructing a new sequence of optimization to outperform standard optimization levels through phase-ordering or optimization selection [33, 5, 16, 17]. This can be done in a coarse [36, 5, 4] or fine-grained manner [14, 45, 16]. End-to-end approaches often provide a framework that can be paired or built with the compiler to offer boosted search or ordering of optimizations [6, 7, 29]. ML-guided approaches build the decision-making capabilities of ML models into single or multiple passes that can act as a standalone ML-enabled pass. Their corresponding ML models can be either built with the compiler using an Ahead-of-Time (AOT) compilation or be used at runtime using an alternative interprocess communication method, e.g., Python named pipes. All things considered, several recent works can be included in multiple classes, therefore, the lists here are not mutually exclusive.
@@
Cummins et al. [16] propose CompilerGym, a fine-grained RL environment to speed up the search for tuning a single or multiple compiler passes defined as action space. The authors leverage OpenAI’s Gym [46] library to provide a flexible environment to tackle the selection and the phase-ordering of optimization. The results showcase the benefits of using their RL agent to reduce the code size of Cbench benchmarks further. Contrary to this work, we propose an ML-guided approach to provide a framework for optimizing passes to be used standalone or inside LLVM’s O3 pipeline. Additionally, our objective is to improve the performance instead of reducing a binary’s code size.
Rotem and Cummins [28] propose a Decision Trees (DS)-based framework to infer branch probabilities for use in providing profile-guided optimizations (PGO). The authors use hand-crafted features for the characterization of code regions and how they are able to reap the benefits of PGO without actually running the program with instrumentation to collect profile data and recompiling. This work is relevant to ours since the approach builds into LLVM and once trained, it can provide PGO with certain profiling information without actually running PGO.
@@
Trofin et al. [15] propose MLGO that uses ML to advise the Inliner as to whether or not a call site should be inlined, in order to minimize the size of the generated code. MLGO framework compiles the ML model into C++ libraries using Ahead-of-Time (AOT) compilation which becomes part of the compiler at build time. Additionally, the authors have extended the code base to provide optimization for the register allocation problem. Later, MLGOPerf [8], extended MLGO to optimize inlining for performance as opposed to code size and showed efficacy on several benchmarks, e.g. SPEC CPU 2006 and Cbench. This work is relevant to us as introduced and upstream the necessary libraries and foundation to train and deploy an ML model to reduce code size reduction.
As mentioned earlier, contrary to MLGO, ACPO focuses on the hard problem of ML performance optimization. We show that ACPO is easily extendible and scalable to any optimization passes and we also provide a modular feature collection class, defined in different scopes of LLVM, Module, Function, CallGraph, and Loops, so users can leverage the ready-to-use set of features in their passes of interest.
III Proposed Methodology
@@
Figure 2 presents the high-level flow of the design of our infrastructure. There are multiple components involved with the ACPO architecture: Autotuner, compiler, ACPOModel abstraction, ML APIs, and ML Framework. For the most part, the autotuner is a wrapper around the compiler and its hooks in the compiler are used to generate model training data. During the inference flow, i.e., when ML models are used for decision-making within the compiler, the autotuner is not used, and we can think of the compiler as being at the top of the stack. The ACPO inference flow begins when an ML-enabled pass is invoked and creates an instance of ACPOModel. The instance of ACPOModel aggregates appropriate features from the analyzed code region and specifies the kind of output required from the ML model. The ACPOModel object then transfers the features and information about the model being invoked to our ML framework via a set of remote procedure call APIs. The APIs provide the ability to send requests, including loading an appropriate model into the ML framework, executing an inference call, and returning the result of the inference back to the compiler. The compiler then applies transformations to the input program as prescribed in the inference output it receives. In the following, we describe in more detail the training and inference flows, and then elaborate on each component of the ACPO framework to show how they fit with the training and inference flows.
@@
An autotuner is a software tool that derives decisions for an optimization scenario (minimization or maximization), by intelligently exploring the design space of compiler optimization parameters [47, 13, 30, 48]. The autotuner parameters can be either coarse- or fine-grained. The former includes the selection and the phase-ordering of different optimization passes as a black box, as well as optimization parameters that apply to the whole translation unit, while the latter facilitates the tuning of optimization parameters for individual code regions of a program, e.g. instructions, loops, functions, etc., independently of each other. In this work, we perform fine-grained autotuning with an autotuner [14], which is built on top of OpenTuner [13]. This allows us to generate meaningful training data and target labels that capture the effect of optimization parameters on code region types of interest.
III-A ACPO Training
A key aspect of creating an ML-driven compiler is the training of models on sample data to ensure that the compiler makes decisions consistently with good results contained in the training data. To do this, we opted for a two-step approach: generate data based on a set of benchmarks the model is expected to reasonably reproduce and then train the model with the collected data. The data generation step is shown in Figure 2a. To generate data for training purposes we rely on the autotuner. The autotuning process allows us to override compiler heuristics with an iterative search for the optimal optimization decisions for a given application. By performing the search on our benchmark programs and recording each step of the search, we can obtain information about the impact of each decision on program performance, and thus collect sufficient information to be able to train the eventual model with a reasonable chance of success. We use the O3 compilation as our baseline and measure impact in terms of speedup relative to the O3 compilation. This process is repeated many times to gather as much data per benchmark as possible to achieve beneficial results on most benchmarks. The resulting trial data is then processed and used as input to the model training.
III-B ACPO Inference